# DeTracker: Motion-decoupled Vehicle Detection and Tracking in Unstabilized Satellite Videos

### v1

@@ V-C1 Effectiveness of GLMD
The GLMD module is designed to decouple satellite platform motion from object motion through global alignment and local refinement across adjacent frames. As shown in Table III, incorporating GLMD into the baseline model leads to a substantial performance gain, improving MOTA by 12.5% (from 40.5% to 52.5%) and IDF1 by 1.1% (from 64.9% to 66.0%).
Fig. 7 provides qualitative evidence of GLMD’s effectiveness. The highlighted purple regions denote larger feature discrepancies relative to the unaligned features. After applying GLMD, the differences between the aligned features and those of the reference frame (frame005) become notably smaller, especially in background areas, and the overall error distribution is more concentrated. The alignment-improvement map further localizes areas where GLMD brings tangible benefits: the dominant light-blue response across most of the scene reflects a consistent global enhancement in feature-space alignment.
@@
To further verify the contribution of individual components within the GLMD module, Table IV presents the ablation results on the SDM-Car-SU dataset under the most challenging setting, U3 (denoted as SDM-Car-SU-3.0). It can be observed that only the global alignment branch leads to an 8.5% improvement in MOTA, indicating that global feature alignment effectively suppresses large-scale drifts caused by platform motion. When a traditional SIFT-based method [20] is introduced, feature-level similarity increases, but the resulting MOTA is lower than that achieved by the learning–based global alignment branch. his suggests that handcrafted keypoint-based methods are less reliable for aligning tiny objects, which are prone to misalignment under severe jitter and low-resolution conditions. By jointly integrating both global alignment and local refienment, the MOTA and IDF1 scores further increase to 52.4% and 63.8%, respectively, demonstrating that local refienment effectively captures small-scale local motions and helps cleanly separate global and local motion patterns.
@@
Figure 8: Feature map visualization results after incorporating different modules. Compared with the baseline, the introduction of the GLMD module enables the network to model global features across the entire image, thereby facilitating subsequent feature alignment. Building upon this, the further integration of the TDFP module enhances the feature representations of tiny objects and extracts richer semantic information.
V-C2 Effectiveness of TDFP
The TDFP module explicitly enhances object features through temporal feature fusion. As shown in Table III, incorporating TDFP increases MT from 681 to 1248 and slightly reduces ML from 312 to 286, indicating a substantial improvement in long-term tracking success. Meanwhile, TDFP leads to notable gains in the primary evaluation metrics, improving MOTA by 17.1% (from 40.5% to 57.6%) and IDF1 by 2.9% (from 64.4% to 67.8%). These gains demonstrate that TDFP enables the tracker to maintain accurate associations over longer periods.
Fig. 8 further illustrates its effect. Compared with the baseline, TDFP produces more discriminative and stable feature representations for small objects, significantly improving cross-frame consistency. As a result, the model remains robust to inter-frame feature fluctuations and preserves reliable object recognition throughout the tracking process.
@@ V-C3 Hyperparameter Analysis
Regarding the loss weighting design, we evaluate two representative settings, $\alpha$ = 0.3, $\beta$ = 0.2 and $\alpha$ = 0.2, $\beta$ = 0.3, which correspond to slightly emphasizing detection or temporal association, respectively. As shown in the table, the best performance is achieved at $\alpha$ = 0.3 and $\beta$ = 0.2, indicating that moderately prioritizing the detection objective leads to more stable object localization, which in turn provides a more reliable foundation for subsequent temporal association and identity preservation.
TABLE V: Comparison on SDM-Car-SU test set with different value of the hyperparameters
MOTA IDF1 IDS FP FN
@@ VI Conclusion
This paper presents DeTracker, a joint detection-and-tracking (JDT) framework tailored for unstabilized satellite videos, aiming to achieve robust multi-object tracking in complex dynamic scenes without relying on any stabilization preprocessing. To explicitly decouple satellite platform motion from true object motion, we introduce the GLMD module, which performs two-stage motion modeling via global alignment and local refinement. This design effectively compensates for large-scale background motion while preserving fine-grained object dynamics, leading to more accurate trajectory estimation and improved motion consistency. In addition, the proposed TDFP module enables multi-scale feature fusion and cross-frame dependency modeling, substantially enhancing the continuity and discriminability of tiny-object representations over time. Furthermore, we release a new simulated benchmark dataset, SDM-Car-SU, which incorporates diverse motion directions and intensities to facilitate systematic evaluation of algorithmic robustness under unstabilized conditions. Extensive experiments on both simulated and real satellite videos demonstrate that DeTracker consistently outperforms state-of-the-art methods, validating its effectiveness for practical satellite video multi-object tracking.
VII Acknowledgments
This work was supported by the National Natural Science Foundation of China (62442107).

### v2

@@ V-C1 Effectiveness of GLMD
The GLMD module is designed to decouple satellite platform motion from object motion through global alignment and local refinement across adjacent frames. As shown in Table III, incorporating GLMD into the baseline model leads to a substantial performance gain, improving MOTA by 12.0% (from 40.5% to 52.5%) and IDF1 by 1.1% (from 64.9% to 66.0%). Notably, the reduction in IDs and the corresponding improvement in IDF1 are primarily attributed to the stable feature alignment and precise target localization capabilities provided by GLMD. By effectively decoupling the complex platform motion from the target intrinsic motion, GLMD significantly mitigates tracking failures and ID switches caused by background drift.
@@
Fig. 8 provides qualitative evidence of GLMD’s effectiveness. The brightly colored regions denote larger feature discrepancies between the current frame (frame006) and the reference frame (frame005). After applying GLMD, the differences between the aligned features and those of the reference frame become notably smaller, especially in background areas, and the overall error distribution is more concentrated. The alignment-improvement map further localizes areas where GLMD brings tangible benefits: the dominant dark-purple response across most of the scene reflects a consistent global enhancement in feature-space alignment.
To further verify the contribution of individual components within the GLMD module, Table IV presents the ablation results on the SDM-Car-SU dataset under the most challenging setting, U3 (denoted as SDM-Car-SU-3). It can be observed that the global alignment branch alone leads to an 8.5% improvement in MOTA, indicating that global feature alignment effectively suppresses large-scale drifts caused by platform motion. When a traditional SIFT-based method [21] is introduced, feature-level similarity increases, but it yields a lower MOTA than the learning-based global alignment branch. This suggests that handcrafted keypoint-based methods are less reliable for aligning tiny objects, which are prone to misalignment under severe jitter and low-resolution conditions. By jointly integrating both global alignment and local refinement, the MOTA and IDF1 scores further increase to 52.4% and 63.8%, respectively, demonstrating that local refinement effectively captures small-scale local motions and helps cleanly separate global and local motion patterns.
V-C2 Effectiveness of TDFP
The TDFP module explicitly enhances object features through temporal feature fusion. As shown in Table III, incorporating TDFP leads to notable gains in the primary evaluation metrics, improving MOTA by 17.1% (from 40.5% to 57.6%) and IDF1 by 2.9% (from 64.9% to 67.8%), indicating a substantial improvement in long-term tracking success. These improvements are attributed to the implicit temporal consistency modeling introduced by the TDFP module at the feature level. By utilizing ConvGRU to recursively propagate and aggregate spatio-temporal features across consecutive frames, the TDFP establishes a cross-frame memory mechanism, thereby maintaining the temporal continuity and discriminative stability of targets within the feature space.
Refer to caption Refer to caption Refer to caption Refer to caption
@@ V-C2 Effectiveness of TDFP
Figure 9: Feature map visualization of adjacent frames after incorporating different modules. Compared with the baseline, the GLMD module effectively corrects background misalignment and reduces static interference (indicated by the white boxes) through global feature alignment. Building on this, the further integration of the TDFP module improves the feature representation of tiny moving objects (indicated by red circles) while filtering out background distractions, such as lane markings (indicated by red boxes).
Fig. 9 further illustrates its effect. Compared with the baseline, TDFP yields more discriminative and stable feature representations for tiny objects, significantly improving cross-frame consistency. As a result, the model remains robust to inter-frame feature fluctuations and preserves reliable object recognition throughout the tracking process.
@@ V-C3 Hyperparameter Analysis
Regarding the loss weight design, we conduct a systematic sensitivity analysis on the classification loss weight $\alpha$ and the bounding-box regression loss weight $\beta$. Given that the overall loss function includes the warp alignment loss, we perform a comprehensive evaluation within the ranges of $\alpha\in\{0.2,0.3\}$ and $\beta\in\{0.1,0.2,0.3\}$ to ensure a balance among the various loss components. As shown in Table V, when $\alpha$ is fixed, a smaller $\beta$ leads to insufficient localization constraints, thereby increasing false positives; conversely, a larger $\beta$ restricts the optimization space for the warp alignment loss, negatively affecting the overall tracking performance. Overall, the optimal performance is achieved when $\alpha=0.3$ and $\beta=0.2$.
TABLE V: Comparison on SDM-Car-SU test set with different values of the hyperparameters
[…]
@@ VI Conclusion
This paper presents DeTracker, a joint-detection-and-tracking (JDT) framework tailored for unstabilized satellite videos, aiming to achieve robust multi-object tracking in complex dynamic scenes without relying on any stabilization preprocessing. To explicitly decouple satellite platform motion from true object motion, we introduce the GLMD module, which performs two-stage motion modeling via global alignment and local refinement. This design effectively compensates for large-scale background motion while preserving fine-grained object dynamics, leading to more accurate trajectory estimation and improved motion consistency. In addition, the proposed TDFP module enables multi-scale feature fusion and cross-frame dependency modeling, substantially enhancing the continuity and discriminability of tiny-object representations over time. Furthermore, we release a new simulated benchmark dataset, SDM-Car-SU, which incorporates diverse motion directions and intensities to facilitate systematic evaluation of algorithmic robustness under unstabilized conditions. Extensive experiments on both simulated and real satellite videos demonstrate that DeTracker consistently outperforms state-of-the-art methods, validating its effectiveness for practical satellite video multi-object tracking.